fields/prob23 7 html x (cm) +173.33 18.33 +100.00 +50.00 [play [pa use step>> reset Please wait for the animation to completely load An electron is shot through a region of constant electric field (time is given in μsec [104 s] and position is given in centimeters Rest a. Which vector most closely shows the direction of the electric field? A. b. What is the magnitude of the electric field?
ughPhyslet Physics by ChrisX+Y gnetism/e fields/prob23.7.html Time: 1.25 y (em) +0.00 +17.50 +0.00 50.00 play pause step> reset Please wait for the animation to completely load. ield (time is given in μsec [10° s] and position is given in centimeters). Restart. on of the electric field? A.
Time: 4.6 vx (cm/micro s)vy +49.00 -16335 +13.35 99.00 play | pause step>> reset Please wait for the animation to completely load. ime is given in usec [10s s] and position is given in centimeters). Restart. he electric field? A.
Physlet Physics by Chris × + v m/e fields/prob23.7.html Time: 5.4 End of Animation x (cm) s) +243.33 53.33 +100.00 50.00 use step>> reset Please wait for the animation to completely load ime is given in μsec [10-6 sl and position is given in centimeters) Restart. he electric field? A.
Answer #1

The electron has final velocity in the direction of B

force on an an electron is opposite to the direction of the electric field as it has -ve charge.

The direction of the field is along F vector.

Force on th electron F = Ee ; e is the field and e is the charge on electron

acceleration of the electron a = F/me = Ee/me ; m - mass of electron

let t is time of flight

displacement s = at2/2 = Eet2/2me

field E = 2sme/et2   

displacement in x = 163.35 cm

time of flight t = 4.6 -1.25 = 3.35 us

E = 2*163.35E-2 * 9.0E-31 /( 1.6E-19 * (3.35E-6)2 )

= 1.6375 N/C

This filed is the x-component which caused acceleration in x-direction

we take file is making -45o with x-axis

actual field = 1.414 * 1.6375 = 2.316 N/C